OpenStudy (anonymous):

What is the period? y=cos(3x)

OpenStudy (anonymous):

@IMStuck maybe you could help me?

OpenStudy (anonymous):

For a periodic function of the form \(\sin bt\) or \(\cos bt\), the period is \(\dfrac{2\pi}{b}\).

OpenStudy (anonymous):

okay and be in this case would be 3?

OpenStudy (anonymous):

That's the value of \(b\), yes.

OpenStudy (imstuck):

Ok, in order to find the period of a function, you have to remember what the original period of the function is. For cosin, the whole thing happens in 2pi. Right? In order to find the period of that (it's a double angle), you divide 2pi by 3 (the 3 is from your (3x). So the new period is 2pi/3.

OpenStudy (anonymous):

okidoki thank you thank you